OpenStudy (anonymous):

What is the change in electric potential difference when +30.0 J of work is done on a charge of +0.010 C? -300 V , +300 V , +3000 V , or -3000 V

OpenStudy (michele_laino):

positive charge will move towards higher potential, so the requested differential of potential will be positive, furthermore, the subsequent relationship, holds: \[\Large \Delta V = \frac{L}{q}\]

OpenStudy (anonymous):

ok:)

OpenStudy (michele_laino):

namely: \[\Large \Delta V = \frac{L}{q} = \frac{{30}}{{{{10}^{ - 2}}}}\]

OpenStudy (anonymous):

i get this? 3000? so my answer is +3000 V ?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay! :) thanks!!

OpenStudy (michele_laino):

:)